# Optimal Allocations under Strongly Pigou-Dalton Criteria: Hidden Layer Structure & Efficient Combinatorial Approach

### v1

@@ 3.1. Layer partition of agents and items
Remark 0.
The Chebyshev distance between different optimal allocations remains to be at most 1 even for the scenario of IND-SUB, as stated in Theorem 22 at later part of this paper. The generalized result Theorem 22 is proved in Section 5, and the reader can find a proof of Theorem 9 (which is much easier) in Appendix B.
As the income $h_{i}$ under different stable allocations $\chi$ does not differ too much for every agent $i$, any solution $\chi$ seems to be acceptable for all of them. However, many questions regarding stable allocations remain to be settled. For example:
@@
    Can we obtain the range of $h_{i}$ under stable allocations?
@@
    How do we count the number of stable allocations?
@@
We introduce a “layer partition” of agents and items in the following, which sheds light on the structure of $\mathcal{S}$ and helps answering the above questions.
@@
As a corollary of Theorem 9, in any stable allocation, each agent $i$ falls into two cases: (1) its income $h_{i}$ is a constant $d$ (for all $\chi\in\mathcal{S}$); or (2) its income $h_{i}$ can be $d$ or d1d-1 for some integer $d>0$. Denote
$\mathsf{layer}_{d}$ $=$ $\{i\mid i\text{ falls into the first case}\}$ (1)
$\mathsf{layer}^{-}_{d}$ $=$ $\{i\mid i\text{ falls into the second case}\}.$ (2)
Obviously, $\mathsf{layer}_{0},\mathsf{layer}_{1},\ldots,\mathsf{layer}^{-}_{1},\mathsf{layer}^{-}_{2},\ldots$ form a partition of agents $[n]$, called the partition of agents. Correspondingly, there also exists a partition of items, denoted by $\mathsf{Layer}_{0}$, $\mathsf{Layer}_{1}$, …, $\mathsf{Layer}^{-}_{1}$, $\mathsf{Layer}^{-}_{2}$, …(definition given below), called the partition of items, and the two layer partitions have the following connection:
Under all stable allocations, the items of some layer always belong to those agents in the same layer (to be clear, throughout this paper, $\mathsf{layer}_{d},\mathsf{Layer}_{d}$ are regarded as in the same layer, whereas $\mathsf{layer}^{-}_{d},\mathsf{Layer}^{-}_{d}$ are regarded as in another layer).
The upcoming lemmas demonstrate how do we explicitly define $\mathsf{layer}_{d}$, $\mathsf{layer}^{-}_{d}$, $\mathsf{Layer}_{d}$, $\mathsf{Layer}^{-}_{d}$, (according to which we can compute the entire partition) and also verify the above connection.
Find a fixed stable allocation $\chi$ (using Theorem 7 Claim 2). Recall the transfers on $\chi$ in Definition 5. See Figure 1. Partition $\{i\mid h_{i}=d\}$ into
$p_{d}$ $=$ $\{i\mid h_{i}=d\text{ and }\exists i\rightarrow j\text{ with }h_{j}=d-1\};$ (3)
$q_{d}$ $=$ $\{i\mid h_{i}=d\text{ and }\exists k\rightarrow i\text{ with }h_{k}=d+1\};$ (4)
$r_{d}$ $=$ $\{i\mid h_{i}=d,i\notin p_{d},\text{ and }i\notin q_{d}\}.$ (5)
Figure 1. Computing layer partition of items from $\chi$.
@@
Note that $p_{d}$ is disjoint with $q_{d}$. Otherwise, there is $k\rightarrow j$ with $h_{k}=h_{j}+2$, contradicting the fact that $\chi$ is stable.
For convenience, denote $s_{d}=p_{d}\cup q_{d-1}$. Moreover, denote by $R_{d}$ the set of items allocated to $r_{d}$ under $\chi$, and $S_{d}$ the set of items allocated to $s_{d}$ under $\chi$.
Lemma 11.
Regardless of which stable allocation we choose, those items in $R_{d}$ are always allocated to $r_{d}$, and those items in $S_{d}$ are always allocated to $s_{d}$.
Proof.
For convenience, let sets $r_{0},s_{1},r_{1},s_{2},\ldots$ be indexed increasingly from 0. So as $R_{0},S_{1},R_{1},S_{2},\ldots$. Moreover, assume that the elements in $r_{0},R_{0}$ have index 0; in $s_{1},S_{1}$ have index $1$; in $r_{1},R_{1}$ have index $2$; and etc.
First, observe that (1) a higher-indexed item cannot be allocated to a lower-indexed agent in any (stable) allocation. This follows from two facts (which can be deduced from the definitions of $p_{d},q_{d},r_{d}$ and the assumption that $\chi$ is stable)
Fact 1. there are no edges from $s_{d}$ to $r_{d-1},s_{d-1},r_{d-2},\ldots$, and
Fact 2. there are no edges from $r_{d}$ to $s_{d},r_{d-1},s_{d-1},\ldots$,
@@
It remains to prove that (2) a lower-indexed item cannot be allocated to a higher-indexed agent in any (stable) allocation. We prove this by induction. Here, for the simplicity of presentation, assume $\max\{h_{i}\}=3$ as shown in Figure 1. First, the items indexed lower than $r_{3}$ cannot be allocated to $r_{3}$ in any stable allocation $\chi^{\prime}$. Otherwise, as the items in $R_{3}$ must be allocated to $r_{3}$ (by the analysis above), some agent in $r_{3}$ receives more than 3 items by the pigeonhole principle, which implies that $\mathbf{p}(\chi^{\prime})$ is non-optimal under $\mathsf{LexiMax}$, meaning that $\chi^{\prime}$ is not stable (Theorem 7). By the same argument, (in any stable allocation $\chi^{\prime}$) the items indexed lower than $s_{3}$ cannot be allocated to $s_{3}$ (otherwise $\mathbf{p}(\chi^{\prime})$ contains more agents receiving $3$ items than $\mathbf{p}(\chi)$ which is impossible), the items indexed lower than $r_{2}$ cannot be given to $r_{2}$, so on and so forth. ∎
Lemma 12.
$\mathsf{layer}_{d}=r_{d}$ and $\mathsf{layer}^{-}_{d}=s_{d}$.
Proof.
According to Lemma 11, in every stable allocation, agents $r_{d}$ will receive and only receive $R_{d}$. Note that $|R_{d}|=d|r_{d}|$. Therefore, in every stable allocation, each agent from $r_{d}$ receives exactly $d$ items (otherwise, there must be one agent receiving more than $d$ and one agent receiving less than $d$, which is worse than equally distribution for the criteria $\mathsf{LexiMin}$ and hence nonstable). In other words, the income of an agent from $r_{d}$ is always $d$. Therefore, $r_{d}\subseteq\mathsf{layer}_{d}$.
For each $i\in p_{d}$, we know $h_{i}=d$ and $h_{i}$ can be reduced to d1d-1 (according to (3)), thus $i\in\mathsf{layer}^{-}_{d}$. Therefore $p_{d}\subseteq\mathsf{layer}^{-}_{d}$. For each $i\in q_{d-1}$, we know $h_{i}=d-1$ and $h_{i}$ can be increased to $d$ (according to (4)), thus $i\in\mathsf{layer}^{-}_{d}$. Therefore $q_{d-1}\subseteq\mathsf{layer}^{-}_{d}$. Together, $s_{d}\subseteq\mathsf{layer}^{-}_{d}$.
@@
It follows that $\mathsf{layer}_{d}=r_{d}$ and $\mathsf{layer}^{-}_{d}=s_{d}$. ∎
Combining the last two lemmas, we have $\mathsf{Layer}_{d}=R_{d}$ and $\mathsf{Layer}^{-}_{d}=S_{d}$.
3.2. Some applications of the layer partition – answering $Q_{1},\ldots,Q_{4}$
@@
Answer of $Q_{3}$.:
Counting stable allocations reduces to counting stable allocations within each layer (and then applying the rule of product). However, this counting problem is #P-hard, as the problem of counting stable allocation in $\mathsf{layer}_{1}$ can be reduced to counting perfect matchings which is already #P-hard.
Answer of $Q_{4}$.:
Finding a profile (of a stable allocation) that minimizes a linear function of $h_{1},\ldots,h_{n}$, e.g. $\sum c_{i}h_{i}$, is easy. Modify the network used in the proof of Theorem 7 Claim 2 as follows: Among the edges from $v_{j}$ to $t$, change the cost of the $k$-th one to be $(k-1)\times A+c_{i}$, where $A$ is a large enough constant. Then, the minimum-cost flow still has the minimum $\mathsf{Congestion}$, and it optimizes $\sum c_{i}h_{i}$. For non-linear functions, it is more difficult. Yet the layer partition still helps break down the task.
Remark 0.
Halpern et al. Halpern et al. (2020) implicitly found a structure similar to our layer structures. Specifically, (in the proof of their Theorem 4) given a fractional MNW allocation, they partition the agents into subsets according to the floor of valuations (so in a certain set, the valuation range of each agent is $[x,x+1)$ for some integer $x$). They imply that the agents in each set must be fully allocated to a certain subsets of items for any fractional MNW allocation, and the partition and correspondence between the subsets of agents and the subsets of items also hold for deterministic MNW allocations. We discover our layer structures independently. Our layer structures are more specific, which can be briefly explained through the valuation range of each agent in a certain layer: for the scenario of DIV, compared to their range of $[x,x+1)$ for some integer $x$, the range in our layer structures is some fixed rational number and for the scenario of IND, compared to their range of $[x,x+1]$ for some integer $x$, the range in our layer structures is $[x,x+1]$ for some integer $x$ or some fixed integer (the range of $[x,x]$). Additionally, to compute the layer partition for deterministic MNW allocations, a fractional MNW allocation is necessary through their framework, while we can compute it directly in the scenario of IND (from a deterministic MNW allocation in their term).
4. Divisible items and agents with 0/1-add valuations
@@
Definition 0 (Stable allocations in scenario of DIV).
@@
Be aware that $\chi^{(b)}$ becomes $\chi^{*}$ after applying the arcs in $C_{1},\ldots,C_{a}$. We obtain that $\mathsf{LexiMin}(\chi^{(b)})=\mathsf{LexiMin}(\chi^{*})$. Further since that $\mathsf{LexiMin}(\chi^{*})<\mathsf{LexiMin}(\chi)$, there exists $i\penalty 10000\ (1\leq i\leq b)$ such that $\mathsf{LexiMin}(\chi^{(i)})<\mathsf{LexiMin}(\chi^{(i-1)})$.
It follows that in $\chi^{(i-1)}$, we have $h_{s}>h_{t}$ (where $s,t$ denote $s_{i},t_{i}$ respectively, for short). It further follows that in $\chi^{(0)}=\chi$, we also have $h_{s}>h_{t}$, since $h_{s}$ never increases and $h_{t}$ never decreases in the sequence $\chi^{(0)},\ldots,\chi^{(i-1)}$. Consequently, $\chi$ admits a narrowing transfer (from $s$ to $t$). ∎
Theorem 16.
1. For any SPD criterion, the optimums coincide with $\mathcal{S}^{*}$. 2. We can find a stable allocation in polynomial time.
Just like we prove Theorem 7 Claim 1 using Lemma 6, we can prove Theorem 16 Claim 1 using Lemma 15 (the proof is omitted).
To find a stable allocation in polynomial time, we can use an approach based on linear programming (LP) which is shown in Section 4.2 (thus we prove Theorem 16 Claim 2). Alternatively, we find a purely combinatorial approach (which is more efficient and more interesting) for finding such an allocation. It utilizes the layer partition with several nontrivial ideas. See details in Section 4.1.
Theorem 17.
@@
A proof of Theorem 17 is related to the LP approach mentioned above and is given in Section 4.2. An alternative proof is similar to the proof of Theorem 9 and is omitted.
4.1. Layer partition of agents and items for divisible items & a combinatorial algorithm for finding a stable allocation
In the following, we extend the layer partition given in Section 3.1 to the divisible case and then present the aforementioned combinatorial algorithm.
According to Theorem 17, profile $\mathbf{p}(\chi)$ is unique for $\chi\in\mathcal{S}^{*}$. In other words, the income $h_{i}$ of each agent $i$ is independent of $\chi$, as long as $\chi$ is stable.
@@
Proof.
Fix $\chi$. Let $L_{d}$ be the items allocated to $\mathsf{layer}_{d}$ under $\chi$. An item in $L_{d^{\prime}}$ ($d^{\prime}>d$) cannot be allocated to $\mathsf{layer}_{d}$ otherwise there is a narrowing transfer in $\chi$. An item in $L_{d^{\prime}}$ ($d^{\prime}<d$) cannot be allocated to $\mathsf{layer}_{d}$ otherwise the allocation is not $\mathsf{LexiMax}$ optimal (can be proved by induction as in the proof of Lemma 11). Therefore, those items in $L_{d}$ can only be allocated to $\mathsf{layer}_{d}$ (even for other stable allocations). We thus obtain the first part of this lemma.
2. In any stable allocation, an item cannot be allocated to different layers. Otherwise there is clearly a simple narrowing transfer. ∎
Lemma 18 implies a layer partition of items and agents, where items $\mathsf{Layer}^{*}_{d}$ and agents $\mathsf{layer}^{*}_{d}$ are in the same layer.
Lemma 19.
@@
Those items in higher layer cannot be given to agents in lower layer – there are no such edges as shown in the proof of Lemma 11 (this also holds for the divisible case). Hence there is no narrowing transfer between layers in $\gamma$. Also, there is no narrowing transfer within any layer of $\gamma$ by the construction of $\gamma$. Together, $\gamma$ admits no narrowing transfer and hence is stable. So, $\gamma\in\mathcal{S}^{*}$. ∎
Lemma 20.
@@
1. $h_{i}=d$ for each $i\in\mathsf{layer}_{d}$. 2. $h_{i}\in[d-1,d]$ for $i\in\mathsf{layer}^{-}_{d}$.
Proof.
@@
    For each $i\in[n]$, it holds that $h_{i}=g_{i}$ (apply Theorem 17 with $\gamma,\chi\in\mathcal{S}^{*}$).
@@
For divisible case, we divide each item into $2n^{2}$ identical (unbreakable) pieces with size $\frac{1}{2n^{2}}$, and solve this indivisible case (with $2n^{2}m$ indivisible pieces) using the algorithm in Theorem 7 Claim 2.
@@
4.2. Find a stable allocation for divisible items using LP and a proof of Theorem 17
@@
Applying Theorem 16 Claim 1, finding a stable allocation reduces to finding an allocation with the minimum $\mathsf{LexiMin}$. We claim that
@@
    the latter further reduces to computing the “fair multi-flow” Nace and Doan (2002) in the network below:
Figure 2. Reduction to fair multi-flow problem.
See Figure 2. There are $n$ source nodes $s^{1},\ldots,s^{n}$ in the first layer $S$. The second layer $R$ consists of only one node $r$ as a relay. The third layer $I$ has $m$ nodes, standing for items. The last layer $T$ has $n$ sink nodes $t^{1},\ldots,t^{n}$. For each agent $d$, there is an arc $(s^{d},r)$ of unlimited capacity. For every item $i$ in $I$, there is an arc $(r,i)$ of capacity 1. If item $i$ can be allocated to agent $d$, an arc $(i,t^{d})$ is added with capacity 1.

### v3

@@ 3.1 Layer partition of agents and items
Proof
Decompose $G$ into paths $P_{1},\ldots,P_{b}$ and a few cycles. Denote by $s_{j},t_{j}$ the starting and ending vertices of $P_{j}$, respectively. We assume that $t_{j}\neq s_{k}$ for $j\neq k$ as in Lemma˜1. Moreover, assume that $s_{1}=s_{2}=i$ (due to $h_{i}\geq h^{\prime}_{i}+2$).
Because $\mathbf{p}(\chi^{(b)})=\mathbf{p}(\chi^{\prime})$, we have $\mathsf{LexiMin}(\chi^{b})=\mathsf{LexiMin}(\chi^{\prime})$. We also know $\mathsf{LexiMin}(\chi^{(0)})=\mathsf{LexiMin}(\chi^{\prime})$ since both $\chi^{(0)},\chi^{\prime}$ are stable (Lemma˜1). Together, $\mathsf{LexiMin}(\chi^{(0)})=\mathsf{LexiMin}(\chi^{(b)})$. It further implies that $\mathsf{LexiMin}(\chi^{(0)})=\mathsf{LexiMin}(\chi^{(1)})=\ldots=\mathsf{LexiMin}(\chi^{b})$. Otherwise, there exists $j$ such that $\mathsf{LexiMin}(\chi^{(j)})$ < $\mathsf{LexiMin}(\chi^{(j-1)})$, which means $\chi^{(j-1)}$ admits a narrowing transfer $s_{j}\rightarrow t_{j}$, implying that $\chi^{(0)}$ admits a narrowing transfer $s_{j}\rightarrow t_{j}$.
Since $\mathsf{LexiMin}(\chi^{(0)})=\mathsf{LexiMin}(\chi^{(1)})=\mathsf{LexiMin}(\chi^{(2)})$, we know $s_{1}\rightarrow t_{1}$, i.e. $i\rightarrow t_{1}$ (recall $s_{1}=i$), is a swapping transfer of $\chi^{(0)}$, whereas $s_{2}\rightarrow t_{2}$, i.e. $i\rightarrow t_{2}$ (recall $s_{2}=i$), is a swapping transfer of $\chi^{(1)}$. It follows that $h_{t_{2}}=h_{i}-2$, therefore $\chi^{(0)}=\chi$ admits a narrowing transfer $i\rightarrow t_{2}$.
Remark 3
We will see at later part of this paper (Theorem˜5.1) that for the scenario of IND-SUB, the Chebyshev distance between different optimal allocations remains to be at most 1. The proof of this generalized result is much more involved and is given in Section˜5.
According to Theorem 3.2, the income $h_{i}$ under different stable allocations $\chi$ does not differ too much for every agent $i$, so every solution $\chi$ seems to be acceptable for all of them. However, many questions regarding stable allocations remain to be settled. For example:
@@
    Can we obtain the range of $h_{i}$ over all stable allocations?
@@
    Is it possible to count efficiently the number of stable allocations?
@@
In what follows, we introduce a construct, called “layer partition” of agents and items, with some nice property (see Lemma˜2 below). This construct elucidates the structure of $\mathcal{S}$ and, as we will demonstrate, provides a pathway to answering the four questions above.
@@
First of all, we point out that each agent $i$ falls into two cases: (1) its income $h_{i}$ is always equal to some constant $d$ for all $\chi\in\mathcal{S}$; or (2) its income $h_{i}$ equals $d$ for some $\chi\in\mathcal{S}$ and equals d1d-1 for some other $\chi\in\mathcal{S}$, for some integer $d>0$. This is a corollary of Theorem˜3.2. Denote by $\mathsf{layer}_{d}$ the set of agents $i$ for which $h_{i}$ always equal to $d$ (for $\chi\in\mathcal{S}$), and by $\mathsf{layer}^{-}_{d}$ the set of agents $i$ for which $h_{i}$ sometimes equal $d$ and sometimes equal d1d-1 (for $\chi\in\mathcal{S}$).
Lemma 2
The set of items allocated to $\mathsf{layer}_{d}$ ($d\geq 0$) is invariant under every stable allocation (denote it by $\mathsf{Layer}_{d}$ henceforth), and the set of items allocated to $\mathsf{layer}^{-}_{d}$ (for $d>0$) is invariant as well (denote it by $\mathsf{Layer}^{-}_{d}$ henceforth). Equivalently speaking, corresponding to the partition of agent $\mathsf{layer}_{0}$, $\mathsf{layer}_{1}$, $\mathsf{layer}^{-}_{1}$, $\mathsf{layer}^{-}_{2}$, $\ldots$, there exists a partition of items $\mathsf{Layer}_{0}$, $\mathsf{Layer}_{1}$, $\mathsf{Layer}^{-}_{1}$, $\mathsf{Layer}^{-}_{2}$, $\ldots$, so that under each stable allocation $\chi\in\mathcal{S}$, the items from $\mathsf{Layer}_{d}$ (or $\mathsf{Layer}^{-}_{d}$, respectively) are always allocated to the agents from $\mathsf{layer}_{d}$ (or $\mathsf{layer}^{-}_{d}$, respectively).
Shortly, items in any given layer can only be allocated to agents in that layer. To be clear, we regard $\mathsf{layer}_{d},\mathsf{Layer}_{d}$ as in the same layer and $\mathsf{layer}^{-}_{d},\mathsf{Layer}^{-}_{d}$ as in another layer.
3.2 Proof of Lemma˜2 and the explicit construction of the layer partition
We prove Lemma˜2 and show how to compute $\mathsf{layer}_{d},\mathsf{layer}^{-}_{d},\mathsf{Layer}_{d},\mathsf{Layer}^{-}_{d}$ efficiently in this subsection. To this end, we introduce some additional notation first.
Fix any stable allocation $\chi\in\mathcal{S}$. Recall the transfers on $\chi$ (and the notation $\rightarrow$) in Definition˜3. Those agents receiving $d$ items under $\chi$ can be classified into three groups:
$p_{d}$ $:=$ $\{i\mid h_{i}=d\text{ and }\exists i\rightarrow j\text{ with }h_{j}=d-1\};$ (1)
$q_{d}$ $:=$ $\{i\mid h_{i}=d\text{ and }\exists k\rightarrow i\text{ with }h_{k}=d+1\};$ (2)
$r_{d}$ $:=$ $\{i\mid h_{i}=d,i\notin p_{d},\text{ and }i\notin q_{d}\}.$ (3)
@@ 3.2 Proof of Lemma˜2 and the explicit construction of the layer partition
See Figure˜2. Observe that $p_{d}$ is disjoint with $q_{d}$. Otherwise, there is $k\rightarrow j$ with $h_{k}=h_{j}+2$, contradicting the fact that $\chi$ is stable. Denote $s_{d}=p_{d}\cup q_{d-1}$. Be aware that the quantity $h_{i}$, the groups $p_{d},q_{d},r_{d}$, and the set $s_{d}$ all depends on the selected stable allocation $\chi$.
Figure 2: Illustration of $p_{d},q_{d},r_{d},s_{d}$.
Denote by $R_{d}$ and $S_{d}$ the set of items allocated to $r_{d}$ and $s_{d}$ under $\chi$, respectively.
Moreover, within this subsection, we rank all items and agents as follows.
  • elements in $r_{0},R_{0}$ have rank 0;  elements in $s_{1},S_{1}$ have rank $1$;
  • elements in $r_{1},R_{1}$ have rank $2$;  elements in $s_{2},S_{2}$ have rank $3$; and etc.
Proposition 1
In the aforementioned stable allocation $\chi$,
  1. 1.
    there is no transfer from $s_{d}$ to any lower-ranked agents (namely, $r_{d-1},s_{d-1},\ldots$);
  2. 2.
    there is no transfer from $r_{d}$ to any lower-ranked agents (namely, $s_{d},r_{d-1},\ldots$);
This easily follows from the definitions of $p_{d},q_{d},r_{d}$. Trivial proofs omitted.
In the following, the rank refers to the rank defined above with respect to the fixed $\chi$.
Proposition 2
In any stable allocation $\chi^{\prime}$, it holds that
  1. 1.
    a higher-ranked item cannot be allocated to a lower-ranked agent.
  2. 2.
    a lower-ranked item cannot be allocated to a higher-ranked agent.
  3. 3.
    Together, items in $R_{d}$ are always assigned to $r_{d}$, and items in $S_{d}$ are always assigned to $s_{d}$.
Proof
1. An item $o$ from $S_{d}$ (or $R_{d}$, respectively) is not attractive to any lower-ranked agent, since otherwise $\chi$ admits a transfer from $s_{d}$ (or $r_{d}$, respectively) to a lower-ranked agent (using just one reallocation of $o$), which contradicts Proposition˜1. As a result, an item from $S_{d}$ (or $R_{d}$, respectively) cannot be allocated to any lower-ranked agent in $\chi^{\prime}$.
@@
2. We prove it by induction. In this proof, for the simplicity of presentation, assume $\max\{h_{i}\}=3$; see Figure˜2. First, the items ranked lower than $r_{3}$ cannot be allocated to $r_{3}$ in $\chi^{\prime}$. Otherwise, as the items in $R_{3}$ must be allocated to $r_{3}$ (by the analysis above), some agent in $r_{3}$ receives more than 3 items by the pigeonhole principle, which implies that $\mathbf{p}(\chi^{\prime})$ is non-optimal under $\mathsf{LexiMax}$, and therefore $\chi^{\prime}$ is not stable (Theorem˜3.1). Similarly, the items ranked lower than $s_{3}$ cannot be allocated to $s_{3}$ in $\chi^{\prime}$, otherwise $\mathbf{p}(\chi^{\prime})$ contains more agents receiving $3$ items than $\mathbf{p}(\chi)$ which is impossible. So on and so forth.
Proposition 3
@@
For $i\in r_{d}$, it holds that $h_{i}$ always equal $d$ for all $\chi\in\mathcal{S}$; and for $i\in s_{d}$, it holds that $h_{i}$ can be $d$ or d1d-1 under different $\chi\in\mathcal{S}$. It follows that $\mathsf{layer}_{d}=r_{d}$ and $\mathsf{layer}^{-}_{d}=s_{d}$.
Proof
According to Proposition˜2 (claim 3), in every stable allocation, agents $r_{d}$ will receive and only receive $R_{d}$. Note that $|R_{d}|=d|r_{d}|$. Therefore, in every stable allocation, each agent from $r_{d}$ receives exactly $d$ items (otherwise, there must be one agent receiving more than $d$ and one agent receiving less than $d$, which is worse than equally distribution for the criteria $\mathsf{LexiMin}$ and hence nonstable). Hence, $h_{i}$ always equals $d$ for $i\in r_{d}$.
For each $i\in p_{d}$, we know $h_{i}=d$ and $h_{i}$ can be reduced to d1d-1 (due to (1)) in another stable allocation. For each $i\in q_{d-1}$, we know $h_{i}=d-1$ and $h_{i}$ can be increased to $d$ (due to (2)) in another stable allocation. Therefore $h_{i}$ can be d1d-1 or $d$ for $i\in p_{d}\cup q_{d-1}=s_{d}$.
Lemma˜2 simply follows from all the analysis above. In particular, combining the last two propositions, we obtain that $\mathsf{Layer}_{d}=R_{d}$ and $\mathsf{Layer}^{-}_{d}=S_{d}$.
3.3 Some applications of the layer partition – answering $Q_{1},\ldots,Q_{4}$
@@
Answer of $Q_{3}$.
Counting stable allocations reduces to counting stable allocations within each layer (and then applying the rule of product).
However, this counting problem is #P-hard, as the problem of counting stable allocation in $\mathsf{layer}_{1}$ can be reduced to counting perfect matchings which is already #P-hard.
Answer of $Q_{4}$.
Finding a profile (of a stable allocation) that minimizes a linear function of $h_{1},\ldots,h_{n}$, e.g. $\sum c_{i}h_{i}$, is easy. Modify the network used in the proof of Theorem˜3.1 Claim 2 as follows: Among the edges from $v_{j}$ to $t$, change the cost of the $k$-th one to be $(k-1)\times A+c_{i}$, where $A$ is a large enough constant. Then, the minimum-cost flow still has the minimum $\mathsf{Congestion}$, and it optimizes $\sum c_{i}h_{i}$. For non-linear functions, it is more difficult. Yet the layer partition still helps break down the task.
Remark 4
Halpern et al. [halpern2020fair] implicitly found a structure similar to our layer structures. Specifically, (in the proof of their Theorem 4) given a fractional MNW allocation, they partition the agents into subsets according to the floor of valuations (so in a certain set, the valuation range of each agent is $[x,x+1)$ for some integer $x$). They imply that the agents in each set must be fully allocated to a certain subsets of items for any fractional MNW allocation, and the partition and correspondence between the subsets of agents and the subsets of items also hold for deterministic MNW allocations. We discover our layer structures independently. Our layer structures are more specific, which can be briefly explained through the valuation range of each agent in a certain layer: for the scenario of DIV, compared to their range of $[x,x+1)$ for some integer $x$, the range in our layer structures is some fixed rational number and for the scenario of IND, compared to their range of $[x,x+1]$ for some integer $x$, the range in our layer structures is $[x,x+1]$ for some integer $x$ or some fixed integer (the range of $[x,x]$). Additionally, to compute the layer partition for deterministic MNW allocations, a fractional MNW allocation is necessary through their framework, while we can compute it directly in the scenario of IND (from a deterministic MNW allocation in their term).
4 Divisible items and agents with 0/1-add valuations
@@
Definition 4(Stable allocations in scenario of DIV)
@@
Be aware that $\chi^{(b)}$ becomes $\chi^{*}$ after applying the arcs in $C_{1},\ldots,C_{a}$. We obtain that $\mathsf{LexiMin}(\chi^{(b)})=\mathsf{LexiMin}(\chi^{*})$. Further since that $\mathsf{LexiMin}(\chi^{*})<\mathsf{LexiMin}(\chi)$, there exists $i~(1\leq i\leq b)$ such that $\mathsf{LexiMin}(\chi^{(i)})<\mathsf{LexiMin}(\chi^{(i-1)})$.
It follows that in $\chi^{(i-1)}$, we have $h_{s}>h_{t}$ (where $s,t$ denote $s_{i},t_{i}$ respectively, for short). It further follows that in $\chi^{(0)}=\chi$, we also have $h_{s}>h_{t}$, since $h_{s}$ never increases and $h_{t}$ never decreases in the sequence $\chi^{(0)},\ldots,\chi^{(i-1)}$. Consequently, $\chi$ admits a narrowing transfer (from $s$ to $t$).
Theorem 4.1
1. For any SPD criterion, the optimums coincide with $\mathcal{S}^{*}$.
2. We can find a stable allocation in polynomial time.
Just like we prove Theorem˜3.1 Claim 1 using Lemma˜1, we can prove Theorem˜4.1 Claim 1 using Lemma˜3 (the proof is omitted).
To find a stable allocation in polynomial time, we can use an approach based on linear programming (LP) which is shown in Section˜4.2 (thus we prove Theorem˜4.1 Claim 2). Alternatively, we find a purely combinatorial approach (which is more efficient and more interesting) for finding such an allocation. It utilizes the layer partition with several nontrivial ideas. See details in Section˜4.1.
Theorem 4.2
@@
A proof of Theorem˜4.2 is related to the LP approach mentioned above and is given in Section˜4.2. An alternative proof is similar to the proof of Theorem˜3.2 and is omitted.
4.1 Layer partition of agents and items for divisible items & a combinatorial algorithm for finding a stable allocation
In the following, we extend the layer partition given in Section˜3.1 to the divisible case and then present the aforementioned combinatorial algorithm.
According to Theorem˜4.2, profile $\mathbf{p}(\chi)$ is unique for $\chi\in\mathcal{S}^{*}$. In other words, the income $h_{i}$ of each agent $i$ is independent of $\chi$, as long as $\chi$ is stable.
@@
Proof
Fix $\chi$. Let $L_{d}$ be the items allocated to $\mathsf{layer}_{d}$ under $\chi$. An item in $L_{d^{\prime}}$ ($d^{\prime}>d$) cannot be allocated to $\mathsf{layer}_{d}$ otherwise there is a narrowing transfer in $\chi$. An item in $L_{d^{\prime}}$ ($d^{\prime}<d$) cannot be allocated to $\mathsf{layer}_{d}$ otherwise the allocation is not $\mathsf{LexiMax}$ optimal (can be proved by induction as in the proof of Proposition˜2). Therefore, those items in $L_{d}$ can only be allocated to $\mathsf{layer}_{d}$ (even for other stable allocations). We thus obtain the first part of this lemma.
2. In any stable allocation, an item cannot be allocated to different layers. Otherwise there is clearly a simple narrowing transfer.
Lemma˜4 implies a layer partition of items and agents, where items $\mathsf{Layer}^{*}_{d}$ and agents $\mathsf{layer}^{*}_{d}$ are in the same layer.
Lemma 5
@@
Those items in higher layer cannot be given to agents in lower layer – there are no such edges as shown in the proof of Proposition˜2 (this also holds for the divisible case). Hence there is no narrowing transfer between layers in $\gamma$. Also, there is no narrowing transfer within any layer of $\gamma$ by the construction of $\gamma$. Together, $\gamma$ admits no narrowing transfer and hence is stable. So, $\gamma\in\mathcal{S}^{*}$.
Lemma 6
@@
1. $h_{i}=d$ for each $i\in\mathsf{layer}_{d}$.
2. $h_{i}\in[d-1,d]$ for $i\in\mathsf{layer}^{-}_{d}$.
Proof
@@
    For each $i\in[n]$, it holds that $h_{i}=g_{i}$ (apply Theorem˜4.2 with $\gamma,\chi\in\mathcal{S}^{*}$).
@@
For divisible case, we divide each item into $2n^{2}$ identical (unbreakable) pieces with size $\frac{1}{2n^{2}}$, and solve this indivisible case (with $2n^{2}m$ indivisible pieces) using the algorithm in Theorem˜3.1 Claim 2.
@@
4.2 Find a stable allocation for divisible items using LP and a proof of Theorem 4.2
@@
Applying Theorem˜4.1 Claim 1, finding a stable allocation reduces to finding an allocation with the minimum $\mathsf{LexiMin}$. We claim that
@@
    the latter further reduces to computing the “fair multi-flow” [nace2002polynomial] in the network below:
Figure 3: Reduction to fair multi-flow problem.
See Figure˜3. There are $n$ source nodes $s^{1},\ldots,s^{n}$ in the first layer $S$. The second layer $R$ consists of only one node $r$ as a relay. The third layer $I$ has $m$ nodes, standing for items. The last layer $T$ has $n$ sink nodes $t^{1},\ldots,t^{n}$. For each agent $d$, there is an arc $(s^{d},r)$ of unlimited capacity. For every item $i$ in $I$, there is an arc $(r,i)$ of capacity 1. If item $i$ can be allocated to agent $d$, an arc $(i,t^{d})$ is added with capacity 1.